This is an electrochemistry question.

A nickel plate electrode was submerged in 0.01 M NiSO4 at 25ºC. Determine the current density when the potential of the nickel electrode is -0.038 V. Assume the electrode reaction is controlled by activation overpotential, \alpha = 0.5, the exchange current density i0 = 1*10-5 A/m^2 , and the activity of Ni2+ is equal to its concentration.
Answer #1

Auswer to the problem The potential Ni-electrode = – 0.038 At equilibrium, Elg = 0 Over potential, y = E-Teq = -0.038-0) = -0